# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

### v1

@@ 5.2 Real Application: Hydropower System
We evaluate DoFlow on real-world hydropower time-series data from Argonne National Laboratory. In this system, water drives a turbine that powers a generator and passes through a transformer before reaching the grid, with control systems monitoring the process. Signals such as water flow, vibration, and electric current are recorded from each component, forming a natural as shown in Fig. 11.
@@
Figure 3 illustrates interventional forecasting in the hydropower system during a true power outage. At this point, the turbine signals ($X_{1,t},X_{2,t}$) break down, leading the entire system into a forced outage. Notably, for $X_{3,t}$ and $X_{4,t}$, DoFlow successfully predicts the characteristic “spikes” in the generator signals that follow the turbine failure, demonstrating that our DAG-based approach captures the complex causal relationships across system. Quantitative metrics, together with the anomaly detection visualizations from the second task, are provided in the Appendix E.4 due to page limits.
@@ 5.3 Real Application: Cancer Treatment Outcomes
Our DoFlow can naturally model interventional forecasting for causal treatment effect estimation in time series. We evaluate it on the cancer-treatment benchmark of Bica et al. [5], which contains daily patient-level tumor volumes and administered therapies. At each day $t$, the treatment variables $\mathbf{X}_{t}=\{X_{i,t}\}_{i=1}^{4}$ (chemotherapy assignment, radiotherapy assignment, chemotherapy dosage, and radiotherapy dosage) act as causal parents of the outcome $Y_{t}$ (tumor volume) on the causal DAG. For each test patient, we use the first $55$ days as observational context and perform interventional rollouts for the next $5$ days under ten distinct treatment options, yielding predicted outcomes $\hat{Y}_{t+1:t+5}$. The normalized RMSE between the true $Y_{t+1:t+5}$ and predicted $\hat{Y}_{t+1:t+5}$ is reported in Appendix E.5.
@@
The detailed settings and results are provided in Appendix E.5. Notably, DoFlow achieves substantial improvements over prior baselines on this causal treatment effect task, as summarized in Table 8.
5.4 Results and Comparisons
@@
  • Training Efficiency. DoFlow trains a separate flow per node in the causal DAG, but each network can be much shallower than transformer models. As shown in Tables 3 and 4 (Appendix), its total model size remains modest, and its training and sampling times are comparable to modern baselines.
@@
  • Interventional Performance. We adapted modern baselines for conditional interventional generation, but DoFlow performs notably better: its RNN–flow design jointly encodes causal histories and propagates interventions through recurrent states, while the flow backbone enables coherent decoding.
@@
  • Counterfactual Generation. As discussed in the Introduction, DoFlow is among the early approaches for generative counterfactual time-series on a causal DAG, complementing work on treatment effects, post-hoc counterfactual explanations, and causal discovery. Such counterfactual generation is crucial for modeling component relationships and enabling post-hoc analyses in domains like healthcare and hydropower.
6 Conclusions and Discussions
@@
We introduced DoFlow, a flow-based generative framework for causal time-series forecasting that unifies observational, interventional, and counterfactual queries on DAG-structured systems. DoFlow achieves accurate forecasting across causal queries and offers explicit likelihoods that support anomaly detection. Moreover, we provide a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions. Experiments on synthetic DAGs and real applications show strong observational forecasts and effective causal query generation. This framework lays the foundation for domains such as healthcare, energy, and other areas where counterfactual and interventional forecasting is critical. DoFlow can also lead to a broader line of research on flow-based causal generative modeling. Extending this framework to partially observed, multi-modal, or physics-informed systems could enable scientific digital twins that reason under interventions and uncertainty. Such integration of causal structure and generative modeling represents a step toward a general theory of inference and control in complex dynamical environments.
@@ Appendix A Proofs
$\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$
@@
which induces a corresponding evolution of the sample density governed by the Liouville continuity equation:
$\partial p_{\theta}(x_{t},s)+\nabla\cdot\big(p_{\theta}(x_{t},s)v(x_{t},s;\hat{H}_{t-1})\big)=0.$ (16)
Here, $p_{\theta}(x_{t}(s),s)$ denotes the time-dependent probability density of the sample $x_{t}$ at time $t$.
@@
$\frac{d}{ds}\log p_{\theta}(x_{t}(s),s)$ $=\frac{\nabla p_{\theta}(x_{t}(s),s)\cdot\partial_{s}x_{t}(s)+\partial_{s}p_{\theta}(x_{t},s)}{p_{\theta}(x_{t}(s),s)}$ (17)
$=\frac{\nabla p_{\theta}\cdot\partial_{s}x_{t}(s)+\partial_{s}p_{\theta}}{p_{\theta}}\Big|_{(x_{t}(s),s)}$ (18)
@@
Starting from an initial sample $z_{t}\sim q(\cdot)$ and integrating from $s=1$ to $s=0$, we have:
$\log p_{\theta,X_{t}}\bigl(x_{t}\mid\hat{H}_{t-1}\bigr)=\log q(z_{t})\;+\;\int_{0}^{1}\nabla_{x}\!\cdot\,v\bigl(x_{t}(s),s;\hat{H}_{t-1}\bigr)\,ds.$ (22)
@@
$\log p_{\theta,X_{\tau+1:T}}\!\left(x_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$ (23)
Proposition 13 (Encoded as a function of the exogenous noise $U$) Let Assumption 4.1 hold. Without loss of generality, suppose the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. At each time $t$, the observed variable is generated by the structural causal model $X_{t}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and that the flow encoder produces $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$. Then there exists a continuously differentiable bijection $g:\mathcal{U}\to\mathcal{Z}$, functionally invariant to $(X_{t-},X_{\mathrm{pa},t-})$, such that,
$Z_{t}\;=\;\Phi_{\theta}\bigl(f(X_{t-},X_{\mathrm{pa},t-},U_{t});H_{t-1}\bigr)\;=\;g\bigl(U_{t}\bigr)\quad\text{a.s.}$ (24)
@@
Corollary 4.5 (Counterfactual recovery) Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we apply the intervention $\mathrm{do}\bigl(X_{t-}=\hat{X}_{t-}^{\mathrm{CF}},X_{\mathrm{pa},t-}=\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$, yielding the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$. Then the decoder recovers the true counterfactual at time step $t$ almost surely:
@@
Proof. We establish the result by induction over time steps $t$. Following the definition of the true counterfactual value $X_{t}^{\mathrm{CF}}$ defined in (14), we denote the corresponding true hidden state $H_{t-1}^{\mathrm{CF}}$, which is updated using the true values $X_{t-1}^{\mathrm{CF}}$ and $X_{\mathrm{pa},t-1}^{\mathrm{CF}}$.
Base case. Let $t_{0}\in\mathcal{I}$ be the first step after the context window. By the intervention rule,
@@
$\small\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};\,H_{t-1}^{\mathrm{F}}\bigr)=\Phi_{\theta}\bigr(f(X_{t-},X_{\text{pa},t-},U_{t});H_{t-1}^{\mathrm{F}}\bigr)=g\bigl(U_{t}\bigr)=\Phi_{\theta}\bigr(f(X_{t-}^{\text{CF}},X_{\text{pa},t-}^{\text{CF}},U_{t});H_{t-1}^{\text{CF}}\bigr)=\Phi_{\theta}(X_{t}^{\mathrm{CF}};H_{t-1}^{\text{CF}}),$ (36)
@@
Appendix B Preliminaries on Continuous Normalizing Flows
For better logical flow in the main text, we omit the CNF preliminaries there and include them here instead. This section introduces the fundamentals of CNFs and their flow-matching training on general data types.
B.1 Neural ODE and Continuous Normalizing Flow:
@@
B.2 Flow Matching
Flow Matching (FM) [41] trains continuous normalizing flows without simulating trajectories by regressing the model velocity $v(x(s),s)$ toward a prescribed target field $u(x(s),s)$. The FM objective is
@@ B.2 Flow Matching
Linear Interpolant: Because $u(x,s)$ is generally intractable unless we condition on the starting point $x_{0}$, Conditional Flow Matching (CFM)  [41] was proposed to find a tractable velocity. Given an observed data $x_{0}$ and an endpoint $x_{1}\sim q(\cdot)$ from base distribution, we can choose an analytic interpolation between $x_{0}$ and $x_{1}$, and define a reference path. An information-preserving and simple choice is the linear interpolant [41]:
@@
Appendix C Data Synthesis

### v2

@@ 5.2 Real Application: Hydropower System
We evaluate DoFlow on real-world hydropower time-series data from Argonne National Laboratory. In this system, water drives a turbine that powers a generator and passes through a transformer before reaching the grid, with control systems monitoring the process. Signals such as water flow, vibration, and electric current are recorded from each component, forming a natural DAG as shown in Fig. 10.
@@
Figure 3 illustrates interventional forecasting in the hydropower system during a true power outage. At this point, the turbine signals ($X_{1,t},X_{2,t}$) break down, leading the entire system into a forced outage. Notably, DoFlow successfully predicts the characteristic “spikes” in the generator signals ($X_{4,t}$ and $X_{5,t}$) that follow the turbine failure, demonstrating that our DAG-based approach captures the complex causal relationships across system. Quantitative metrics, together with the anomaly detection visualizations from the second task, are provided in the Appendix F.3.
@@ 5.3 Real Application: Cancer Treatment Outcomes
Our DoFlow can naturally model interventional forecasting for causal treatment effect estimation in time series. The causal effects on time series were thoroughly introduced in Section 1.1. We evaluate DoFlow on the cancer-treatment benchmark of Bica et al. (2020a), which contains daily patient-level tumor volumes (outcome) and administered therapies (actions). The action variables $\mathbf{X}_{t-1}=\{X_{i,t-1}\}_{i=1}^{4}$ (chemotherapy assignment, radiotherapy assignment, chemotherapy dosage, and radiotherapy dosage) act as causal parents of the outcome $Y_{t}$ (tumor volume) on the causal DAG. For each test patient, we use the first $55$ days as observational context and perform interventional rollouts for the next $5$ days under ten distinct treatment options, yielding predicted outcomes $\hat{Y}_{t+1:t+5}$. The normalized RMSE between the true $Y_{t+1:t+5}$ and predicted $\hat{Y}_{t+1:t+5}$ is reported in Appendix F.4.
@@
The detailed settings and results are provided in Appendix F.4. Notably, DoFlow achieves substantial improvements over prior baselines on this causal treatment effect task, as summarized in Table 8.
5.4 Discussions and Limitations
@@
  • Training Efficiency. DoFlow trains a separate flow per node in the causal DAG, but each network can be much shallower than transformer models. As shown in Tables 2 and 3, its total model size remains modest, and its training and sampling times are comparable to modern baselines.
@@
  • Interventions. We adapt modern baselines for conditional intervention generation, but DoFlow performs notably better: its RNN–flow design jointly encodes causal histories and propagates interventions through recurrent states, while the flow backbone enables coherent decoding.
@@
  • Counterfactuals. As discussed in the Introduction, DoFlow is among the early approaches for generating counterfactual time-series on a causal DAG, complementing work on treatment effects and causal discovery on time series. Such counterfactual generation is crucial for modeling component relationships and enabling post-hoc analyses in crucial areas.
  • Limitations - Known DAG. DoFlow currently assumes a known and correctly specified causal DAG, as well as causal sufficiency (see Assumption B.1). In practice, the DAG may come from known physical relationships or from upstream causal discovery methods as discussed in Section 1.1. A natural direction for future work is to couple DoFlow with time-series causal discovery or deconfounding causal generative models (Almodóvar et al., 2025; Xia et al., 2022), allowing the model to operate on learned temporal causal structures for more reliable causal reasoning.
6 Conclusions
@@
We introduced DoFlow, a flow-based generative framework for causal time-series forecasting that unifies observational, interventional, and counterfactual queries on DAG-structured causal systems. Moreover, we provide a supporting counterfactual recovery result for DoFlow (Corollary 4.5) under certain assumptions. Experiments on various synthetic DAGs and real applications show strong observational forecasts and effective prediction under causal queries. This framework lays the foundation for domains such as energy, healthcare, and other areas where counterfactual and interventional forecasting is critical. The integration of causal structure and generative modeling represents a step toward a general theory of inference and control in complex dynamical environments.
Acknowledgment
This work is partially supported by NSF CAREER CCF-1650913, NSF DMS-2134037, CMMI-2015787, CMMI-2112533, DMS-1938106, DMS-1830210, and the Coca-Cola Foundation.
@@ Appendix A Proofs
$\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right],$
where $\hat{x}_{t}=\Phi_{\theta}^{-1}(z_{t};\hat{H}_{t-1}),\ t\in\{\tau+1,\dots,T\}$.
@@
which induces a corresponding evolution of the sample density governed by Liouville’s continuity equation [Chen et al., 2018]:
$\partial p_{\theta}(x_{t},s\mid\hat{H}_{t-1})+\nabla\cdot\big(p_{\theta}(x_{t},s\mid\hat{H}_{t-1})v(x_{t},s;\hat{H}_{t-1})\big)=0.$ (18)
Here, $p_{\theta}(x_{t}(s),s)$ denotes the probability density of the sample $x_{t}(s)$ at ODE-time $s$.
@@
$\frac{d}{ds}\log p_{\theta}(x_{t}(s),s\mid\hat{H}_{t-1})$ $=\frac{\nabla p_{\theta}\cdot\partial_{s}x_{t}(s)+\partial_{s}p_{\theta}}{p_{\theta}}\Big|_{(x_{t}(s),s)}$ (19)
@@
Starting from a base sample $z_{t}\sim q(\cdot)$ and integrating from $s=1$ to $s=0$, we have:
$\log p_{\theta,X_{t}}\bigl(\hat{x}_{t}\mid\hat{H}_{t-1}\bigr)=\log q(z_{t})\;+\;\int_{0}^{1}\nabla_{x}\!\cdot\,v\bigl(x_{t}(s),s;\hat{H}_{t-1}\bigr)\,ds.$ (23)
@@
$\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right],$ (24)
where $\hat{x}_{t}=\Phi_{\theta}^{-1}(z_{t};\hat{H}_{t-1}),\ t\in\{\tau+1,\dots,T\}$. ∎
Proposition 14 (Encoded as a function of the exogenous noise $U$). Let Assumption 4.1 hold. Without loss of generality, suppose the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. At each time $t$, the observed variable is generated by the structural causal model $X_{t}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and that the flow encoder produces $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$. Then there exists a continuously differentiable bijection $g:\mathcal{U}\to\mathcal{Z}$, functionally invariant to $H_{t-1}$, such that,
$Z_{t}\;=\;\Phi_{\theta}\bigl(X_{t};H_{t-1}\bigr)\;=\;\Phi_{\theta}\bigl(f(X_{t-},X_{\mathrm{pa},t-},U_{t});H_{t-1}\bigr)\;=\;g\bigl(U_{t}\bigr)\quad\text{a.s.}$
@@
Corollary 4.5 (Counterfactual recovery). Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we obtain the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$ from the predicted counterfactual past $(\hat{X}_{t-}^{\mathrm{CF}},\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$. The decoder recovers the true counterfactual at time step $t$ almost surely:
@@
Proof. We establish the result by induction over time steps $t$. Following the definition of the true counterfactual value $X_{t}^{\mathrm{CF}}$ defined in (15), we denote the corresponding true hidden state $H_{t-1}^{\mathrm{CF}}$, which is updated using the true counterfactual values $X_{t-1}^{\mathrm{CF}}$ and $X_{\mathrm{pa},t-1}^{\mathrm{CF}}$.
Base case. Let $t_{0}\in\mathcal{I}$ be the first interventional step after the context window. By the intervention rule,
@@
$\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};\,H_{t-1}^{\mathrm{F}}\bigr)=\Phi_{\theta}\bigr(f(X_{t-},X_{\text{pa},t-},U_{t});H_{t-1}^{\mathrm{F}}\bigr)=g\bigl(U_{t}\bigr)=\Phi_{\theta}\bigr(f(X_{t-}^{\text{CF}},X_{\text{pa},t-}^{\text{CF}},U_{t});H_{t-1}^{\text{CF}}\bigr)=\Phi_{\theta}(X_{t}^{\mathrm{CF}};H_{t-1}^{\text{CF}}),$ (36)
@@
Appendix B Causal SCM Assumptions and Theoretical Comparisons to BGM
B.1 Causal SCM Assumptions
Let $\mathbf{X}_{\tau+1:T}^{(\gamma_{\mathcal{I}})}$ denote the potential trajectory under the intervention $\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})$, where $\mathcal{I}\subseteq[K]\times\{\tau{+}1,\dots,T\}$. We assume the following standard conditions for the structural causal model on the given DAG:
Assumption B.1.
(1) Unconfoundedness. The exogenous noises $(U_{i,t})$ are mutually independent across $(i,t)$, and all common causes of observed nodes are included in the DAG, i.e., no unobserved confounders for the interventions considered.
(2) Consistency. If the system is run under $\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})$, then the observed trajectory equals the corresponding potential trajectory under the regime actually applied: $\mathbf{X}_{\tau+1:T}=\mathbf{X}_{\tau+1:T}^{(\gamma_{\mathcal{I}})}$.
(3) Positivity. For any $(x_{i,t-},x_{\mathrm{pa}(i),t-})$ in the support of $(X_{i,t-},X_{\mathrm{pa}(i),t-})$, the conditional density of $X_{i,t}$ given these parents is positive in a neighborhood of the intervention value $\gamma_{i,t}$:
$p\bigl(X_{i,t}=x\,\big|\,X_{i,t-}=x_{i,t-},X_{\mathrm{pa}(i),t-}=x_{\mathrm{pa}(i),t-}\bigr)>0.$
(4) No Interference. For any two intervention schedules $\mathcal{I}$ and $\mathcal{J}$ with the same assigned values on a given index set, the corresponding potential trajectories coincide on that set. Interventions affect other nodes only through the edges of the DAG.
Remark B.2.
Assumption B.1 (1) imposes causal sufficiency: all common causes of the observed nodes are included in the DAG, so the exogenous noises $(U_{i,t})$ are mutually independent across $(i,t)$. This rules out time–varying latent confounders that simultaneously influence multiple series. In settings where unobserved confounding is suspected, our identification and recovery guarantees for interventional and counterfactual distributions need not hold.
Extending DoFlow beyond the causal sufficiency assumption to settings with latent confounders is an important direction for future work. Recent causal generative frameworks such as DeCaFlow [Almodóvar et al., 2025] and Neural Causal Models [Xia et al., 2022] explicitly address hidden confounding problems. DeCaFlow augments causal normalizing flows with latent confounders and proxy variables to recover identifiable interventional and counterfactual queries. Besides, Neural Causal Models use neural SCMs with correlated exogenous variables and graphical constraints to decide and estimate identifiable counterfactuals from observational data. A promising direction is to combine such deconfounding mechanisms with our time-indexed CNF parameterization so that DoFlow can handle deconfounded temporal DAG.
B.2 Theoretical Comparisons to BGM
Bijective Generation Mechanisms (BGM). The BGM framework [Nasr-Esfahany et al., 2023] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
$\hat{f}(X_{t-},X_{\text{pa},t-},U_{t})$ $\stackrel{{\scriptstyle d}}{{=}}\ f(X_{t-},X_{\text{pa},t-},U_{t})$
Relationships. Because of Proposition 4.3 (enabled by the CNF), DoFlow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U_{t}$. However, they are not the same: BGM additionally requires observational distribution matching (as formalized in (A4)) to obtain class-level identifiability, whereas our proofs of Proposition 4.3 and Corollary 4.5 do not assume (A4) and instead rely on the model-specific condition (A3).
Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $g:U_{t}\to Z_{t}$ (independent of $(X_{t-},X_{\text{pa},t-}$), so the induced mechanism $\hat{f}$ implemented by DoFlow is bijective in the exogenous noise, i.e., DoFlow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [Nasr-Esfahany et al., 2023].
Appendix C Preliminaries on Continuous Normalizing Flows
For better logical flow in the main text, we move the CNF preliminaries here. Continuous normalizing flows (CNFs) have been successful for statistical sampling [Tian et al., 2024, Wu and Xie, 2025] and counterfactual inference [Wu et al., 2025, Chao et al., 2024]. This appendix section reviews CNF fundamentals and flow-matching training for general data types.
C.1 Neural ODE and Continuous Normalizing Flow:
@@
C.2 Flow Matching
Flow Matching (FM) [Lipman et al., 2023] trains continuous normalizing flows without simulating trajectories by regressing the model velocity $v(x(s),s)$ toward a prescribed target field $u(x(s),s)$. The FM objective is
@@ C.2 Flow Matching
Linear Interpolant: Because $u(x,s)$ is generally intractable unless we condition on the starting point $x_{0}$, Conditional Flow Matching (CFM)  [Lipman et al., 2023] was proposed to find a tractable velocity. Given an observed data $x_{0}$ and an endpoint $x_{1}\sim q(\cdot)$ from base distribution, we can choose an analytic interpolation between $x_{0}$ and $x_{1}$, and define a reference path. An information-preserving and simple choice is the linear interpolant [Lipman et al., 2023]:
@@
Appendix D Data Synthesis and Metrics Definition